Show that the equation z 3 -1- 2e‘ = 0 has exactly one solution. State clearly every result you use and show that the conditions are satisfied. (Hint: Start by showing that there is at least one solution and then exclude the possibility of more solutions.)
Answer #1

Criven: 23_1-2e=0 Then het foz) = 23-1-20 fco)=– 1-2e40 (1) fer) - 1-1-ze co 7- ze = 27-1-2e do f(3) [3ceca there is som Tus